# Tetrahedron Conjecture in the $\ell_{2}$-norm

### v1

@@ 1 Introduction
The famous Tetrahedron Conjecture of Turán [Tur41] asserts that $\mathrm{ex}(n,K_{4}^{3})=|\mathbb{C}_{n}|$. Various distinct constructions achieving the same bound as $\mathbb{C}_{n}$ were discovered later (see, e.g., [Bro83, Kos82, FDF88, Fro08]). In particular, this implies that the Turán problem for $K_{4}^{3}$ does not exhibit the stability property (see [LM22, Proposition 1.7]), which is likely one reason for the notorious difficulty of the conjecture. Successively better upper bounds for $\mathrm{ex}(n,K_{4}^{3})$ have been obtained by many researchers (see e.g., [dC88, CL99, Raz10a, BT11]), with the current record established by the computer-assisted flag algebra framework of Razborov [Raz07].
@@
The construction $\mathbb{C}_{n}$ shows that $\pi_{\ell_{2}}(K_{4}^{3})\geq\frac{1}{3}$. Using Razborov’s flag algebra framework, Balogh–Clemen–Lidický [BCL22] proved, with computer assistance, that $\pi_{\ell_{2}}(K_{4}^{3})\leq\frac{1}{3}$ and thus established the following theorem.
Theorem 1.1 ([BCL22, Theorem 1.2]).
@@
They further conjectured (see [BCL22, Conjecture 1.3]) that for sufficiently large $n$, the construction $\mathbb{C}_{n}$ is the unique extremal configuration for the Turán problem of $K_{4}^{3}$ in the $\ell_{2}$-norm, in contrast to the $\ell_{1}$-norm setting, where exponentially many extremal constructions may exist. Our main result confirms this conjecture.
@@ 2 Mantel’s theorem in vertex-colored graphs
In this section, we present the proof of Theorem 2.3, derived by computer using the flag algebra method of Razborov [Raz07], which is also described in e.g. [Raz10b, BaberTalbot11, dCSdOFS16, GGH+22]. Since this method is well-known by now, we will be very brief. In particular, we omit many definitions, referring the reader to [Raz07, Raz10b] for any missing notions. Roughly speaking, a flag algebra proof using 0-flags on $m$ vertices of an upper bound $u\in\mathbb{R}$ on the given objective function $f$ consists of an identity
@@
which is asymptotically true for any admissible $\mathcal{H}$ with $|V(\mathcal{H})|\to\infty$, where the $\mathrm{SOS}$-term can be represented as a sum of squares (as described e.g. in [Raz10b, Section 3]), each coefficient $c_{F}\in\mathbb{R}$ is non-negative, and $\mathcal{F}_{m}^{0}$ consists of isomorphism types of 0-flags (unlabeled $3$-graphs) with $m$ vertices. If $f(\mathcal{H})$ can be represented as a linear combination of the densities of members of $\mathcal{F}_{m}^{0}$ in $\mathcal{H}$ then finding the smallest possible $u$ amounts to solving a semi-definite program (SDP) with $|\mathcal{F}_{m}^{0}|$ linear constraints (so we write the size of $\mathcal{F}_{m}^{0}$ in each case to give the reader some idea of the size of the programs that we had to solve).
@@ 3.1 Preliminaries
The following statements can be verified by straightforward calculations.
@@ 4 Concluding remarks
The construction $\mathbb{B}_{n}$ does not contain the complete $3$-graph on five vertices $K_{5}^{3}$. It follows that $\pi_{\ell_{2}}(K_{5}^{3})\geq\frac{5}{8}$. Using computer-assisted flag algebra computations, Balogh–Clemen–Lidický [BCL22] proved that $\pi_{\ell_{2}}(K_{5}^{3})\leq\frac{5}{8}$, and thus, $\pi_{\ell_{2}}(K_{5}^{3})=\frac{5}{8}$. They further conjectured (see [BCL22, Conjecture 1.4]) that for large $n$, the construction $\mathbb{B}_{n}$ is the unique extremal construction for the Turán problem of $K_{5}^{3}$ in the $\ell_{2}$-norm. We will confirm this conjecture along with a special case of a related conjecture of Frankl–Gryaznov–Talebanfard [FGT22] in a forthcoming work, after establishing a Turán theorem in vertex-colored graphs.

### v2

@@ 1 Introduction
The famous Tetrahedron Conjecture of Turán [Tur41] asserts that $\mathrm{ex}(n,K_{4}^{3})=|\mathbb{C}_{n}|$. Various distinct constructions achieving the same bound as $\mathbb{C}_{n}$ were discovered later (see, e.g., [Bro83, Kos82, FDF88, Fro08]). In particular, this implies that the Turán problem for $K_{4}^{3}$ does not exhibit the stability property (see [LM22, Proposition 1.7]), which is likely one reason for the notorious difficulty of the conjecture. Successively better upper bounds for $\mathrm{ex}(n,K_{4}^{3})$ have been obtained by many researchers (see e.g., [dC88, CL99, Raz10, BT11]), with the current record established by the computer-assisted flag algebra framework of Razborov [Raz07].
@@
Turán problems under the $\ell_{2}$-norm were systematically studied by Balogh–Clemen–Lidický in [BCL22a], and later extended to the more general $(t,p)$-norms as well as to degenerate hypergraphs in [CIL+24, GLMP25].
The construction $\mathbb{C}_{n}$ shows that $\pi_{\ell_{2}}(K_{4}^{3})\geq\frac{1}{3}$. Using Razborov’s flag algebra framework, Balogh–Clemen–Lidický [BCL22b] proved, with computer assistance, that $\pi_{\ell_{2}}(K_{4}^{3})\leq\frac{1}{3}$ and thus established the following theorem.
Theorem 1.1 ([BCL22b, Theorem 1.2]).
@@
They further conjectured (see [BCL22b, Conjecture 1.3]) that for sufficiently large $n$, the construction $\mathbb{C}_{n}$ is the unique extremal configuration for the Turán problem of $K_{4}^{3}$ in the $\ell_{2}$-norm, in contrast to the $\ell_{1}$-norm setting, where exponentially many extremal constructions may exist. Our main result confirms this conjecture.
@@ 2 Mantel’s theorem in vertex-colored graphs
In this section, we present the proof of Theorem 2.3, derived by computer using the flag algebra method of Razborov [Raz07], which is also described in e.g. [Raz10, BT11, dCSdOFS16, GGH+22]. Since this method is well-known by now, we will be very brief. In particular, we omit many definitions, referring the reader to [Raz07, Raz10] for any missing notions. Roughly speaking, a flag algebra proof using 0-flags on $m$ vertices of an upper bound $u\in\mathbb{R}$ on the given objective function $f$ consists of an identity
@@
which is asymptotically true for any admissible $\mathcal{H}$ with $|V(\mathcal{H})|\to\infty$, where the $\mathrm{SOS}$-term can be represented as a sum of squares (as described e.g. in [Raz10, Section 3]), each coefficient $c_{F}\in\mathbb{R}$ is non-negative, and $\mathcal{F}_{m}^{0}$ consists of isomorphism types of 0-flags (unlabeled $3$-graphs) with $m$ vertices. If $f(\mathcal{H})$ can be represented as a linear combination of the densities of members of $\mathcal{F}_{m}^{0}$ in $\mathcal{H}$ then finding the smallest possible $u$ amounts to solving a semi-definite program (SDP) with $|\mathcal{F}_{m}^{0}|$ linear constraints (so we write the size of $\mathcal{F}_{m}^{0}$ in each case to give the reader some idea of the size of the programs that we had to solve).
@@ 3.1 Preliminaries
The following statements can be verified by straightforward calculations (a proof of Fact 3.1 is provided in the appendix of the arXiv version).
@@ 4 Concluding remarks
The construction $\mathbb{B}_{n}$ does not contain the complete $3$-graph on five vertices $K_{5}^{3}$. It follows that $\pi_{\ell_{2}}(K_{5}^{3})\geq\frac{5}{8}$. Using computer-assisted flag algebra computations, Balogh–Clemen–Lidický [BCL22b] proved that $\pi_{\ell_{2}}(K_{5}^{3})\leq\frac{5}{8}$, and thus, $\pi_{\ell_{2}}(K_{5}^{3})=\frac{5}{8}$. They further conjectured (see [BCL22b, Conjecture 1.4]) that for large $n$, the construction $\mathbb{B}_{n}$ is the unique extremal construction for the Turán problem of $K_{5}^{3}$ in the $\ell_{2}$-norm. We will confirm this conjecture along with a special case of a related conjecture of Frankl–Gryaznov–Talebanfard [FGT22] in a forthcoming work, after establishing a Turán theorem in vertex-colored graphs.
@@ Acknowledgments
Appendix
Proof of Fact 3.1.
Let $n_{i}\coloneqq|V_{i}|$ for $i\in[3]$. It is enough to show that $|n_{i}-n_{j}|\leq 1$ for all $i,j\in[3]$. Suppose on the contrary that this is false. We will construct a $3$-graph $\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]$ with $V_{1}^{\prime}\cup V_{2}^{\prime}\cup V_{3}^{\prime}=[n]$ such that $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}>\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$, thus proving the lemma. Without loss of generality, we may assume that $n_{1}=\max\{n_{1},n_{2},n_{3}\}$. Since $n_{1}+n_{2}+n_{3}=n\geq 6$, we have $n_{1}\geq 2$ by average. By symmetry, there are two cases.
Case 1: $n_{1}\geq n_{2}\geq n_{3}$ and $n_{1}\geq n_{3}+2$.
Let $|V_{1}^{\prime}|=n_{1}-1$, $|V_{2}^{\prime}|=n_{2}$, and $|V_{3}^{\prime}|=n_{3}+1$. Some calculations show that
$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=$ $n_{1}^{2}n_{3}+2n_{1}(n_{2}^{2}+n_{3}^{2}-n_{2}-n_{3})+n_{2}^{3}-n_{3}^{3}-2n_{2}^{2}n_{3}$
$-3n_{2}n_{3}^{2}-\frac{7}{2}n_{2}^{2}-\frac{1}{2}n_{3}^{2}+2n_{2}n_{3}+\frac{5}{2}n_{2}+\frac{1}{2}n_{3}.$ (33)
It suffices to show that this expression is strictly positive. When $n_{3}=0$, this is a non-decreasing linear function of $n_{1}$. When $n_{3}\geq 1$, this is a quadratic function of $n_{1}$, and it is increasing in $n_{1}\geq 2$. So the minimum value of (Appendix) is attained when $n_{1}=\max\{n_{2},n_{3}+2\}$.
If $n_{1}=n_{2}$, then it follows from (Appendix) that
$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=3n_{1}^{3}-\frac{11}{2}n_{1}^{2}-n_{3}n_{1}^{2}+\frac{5}{2}n_{1}-n_{3}^{2}n_{1}-n_{3}^{3}-\frac{1}{2}n_{3}^{2}+\frac{1}{2}n_{3},$
which is increasing in $n_{1}\geq n_{3}+2$. So it is enough to assume that $n_{1}=n_{2}=n_{3}+2$. Then we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=6n_{3}^{2}+13n_{3}+7>0$.
If $n_{1}=n_{3}+2$, then $n_{2}$ is one of $n_{3}+2,n_{3}+1$ and $n_{3}$. Substituting it into (Appendix), then $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$ is $6n_{3}^{2}+13n_{3}+7$, $6n_{3}^{2}+3n_{3}$ and $6n_{3}^{2}-n_{3}$, respectively. The assumptions $n_{1}=n_{3}+2$ and $n\geq 6$ imply that $n_{3}\geq 1$. Therefore, we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}>0$, as desired.
Case 2: $n_{1}\geq n_{3}\geq n_{2}$ and $n_{1}\geq n_{2}+2$.
Let $|V_{1}^{\prime}|=n_{1}-1$, $|V_{2}^{\prime}|=n_{2}+1$, and $|V_{3}^{\prime}|=n_{3}$. Some calculations show that
$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=$ $\left(1-n_{1}\right)n_{2}^{2}-2n_{2}\left(n_{1}^{2}+n_{3}^{2}-2n_{1}-n_{3}+1\right)+n_{1}^{3}-\frac{9}{2}n_{1}^{2}$
$+\frac{13}{2}n_{1}-n_{3}^{3}-\frac{1}{2}n_{3}^{2}+\frac{9}{2}n_{3}+3n_{1}^{2}n_{3}+2n_{1}n_{3}^{2}-8n_{1}n_{3}-3.$ (34)
This is a quadratic function of $n_{2}$, and it is decreasing in $n_{2}\geq 0$. So it is enough to consider the case $n_{2}=\min\{n_{3},n_{1}-2\}$.
If $n_{2}=n_{3}$, then it follows from (Appendix) that
$\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}\\ =n_{1}^{3}+n_{2}n_{1}^{2}-\frac{9}{2}n_{1}^{2}+n_{2}^{2}n_{1}+\frac{13}{2}n_{1}-4n_{1}n_{2}-3n_{2}^{3}+\frac{5}{2}n_{2}^{2}+\frac{5}{2}n_{2}-3,$ (35)
which is increasing in $n_{1}\geq n_{2}+2$. So it is enough to assume that $n_{1}=n_{2}+2$ and $n_{3}=n_{2}$. Then we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}=6n_{2}^{2}-n_{2}>0$.
If $n_{2}=n_{1}-2$, then $n_{3}$ is one of $n_{1}$, $n_{1}-1$ and $n_{1}-2$. Substituting it into (Appendix), then $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}$ is $6n_{1}^{2}-11n_{1}+5$, $6n_{1}^{2}-15n_{1}+9$ and $6n_{1}^{2}-25n_{1}+26$, respectively. The assumptions $n_{1}\geq n_{2}+2$ and $n\geq 6$ imply that $n_{1}\geq 3$. Therefore, we have $\lVert\mathbb{C}[V_{1}^{\prime},V_{2}^{\prime},V_{3}^{\prime}]\rVert_{2}-\lVert\mathbb{C}[V_{1},V_{2},V_{3}]\rVert_{2}>0$, as desired.
Authors
Levente Bodnár
Mathematics Institute and DIMAP
University of Warwick
Coventry, CV4 7AL, UK
bodnalev@gmail.com
Wanfang Chen
School of Mathematical Sciences
University of Science and Technology of China
Hefei, 230026, China
a372959313@gmail.com
Jinghua Deng
Center for Discrete Mathematics
Fuzhou University
Fujian, 350003, China
Jinghua_deng@163.com
Jianfeng Hou
Center for Discrete Mathematics
Fuzhou University
Fujian, 350003, China
jfhou@fzu.edu.cn
Xizhi Liu
School of Mathematical Sciences
University of Science and Technology of China
Hefei, 230026, China
liuxizhi@ustc.edu.cn
Jialei Song
School of Mathematical Sciences
Key Laboratory of MEA (Ministry of Education) & Shanghai Key Laboratory of PMMP
East China Normal University
Shanghai 200241, China
jialsong@foxmail.com
Jiabao Yang
School of Mathematics
Nanjing University
Nanjing, 210093, China
jbyang1215@nju.edu.cn
Yixiao Zhang
Center for Discrete Mathematics
Fuzhou University
Fujian, 350003, China
fzuzyx@gmail.com